# Particle Semi-Implicit Variational Inference

### v1

@@ Abstract
Semi-implicit variational inference (SIVI) enriches the expressiveness of variational families by utilizing a kernel and a mixing distribution to hierarchically define the variational distribution. Existing SIVI methods parameterize the mixing distribution using implicit distributions, leading to intractable variational densities. As a result, directly maximizing the evidence lower bound (ELBO) is not possible and so, they resort to either: optimizing bounds on the ELBO, employing costly inner-loop Markov chain Monte Carlo runs, or solving minimax objectives. In this paper, we propose a novel method for SIVI called Particle Variational Inference (PVI) which employs empirical measures to approximate the optimal mixing distributions characterized as the minimizer of a natural free energy functional via a particle approximation of an Euclidean–Wasserstein gradient flow. This approach means that, unlike prior works, PVI can directly optimize the ELBO; furthermore, it makes no parametric assumption about the mixing distribution. Our empirical results demonstrate that PVI performs favourably against other SIVI methods across various tasks. Moreover, we provide a theoretical analysis of the behaviour of the gradient flow of a related free energy functional: establishing the existence and uniqueness of solutions as well as propagation of chaos results.
@@ 1 Introduction
In Bayesian inference, a quantity of vital importance is the posterior $p(x|y)={p(\cdot,y)}/{\int p(x,y)\mathrm{d}x}$ where $p(x,y)$ is a probabilistic model; $y$ denotes the observed data and $x$ the model parameters. Here and throughout we assume that the distributions and kernels of interested admit densities. An ever-present issue in Bayesian inference is that the posterior is often intractable as the normalizing constant is available only in the form of an integral and requires methods for approximating it. One popular method is variational inference (VI) (Jordan, 1999; Wainwright et al., 2008; Blei et al., 2017). The essence of VI is to approximate the posterior with a member from a variational family $\mathcal{Q}$ where each element of $\mathcal{Q}$ is a distribution $q_{\theta}(x)$ (called “variational distribution”) parameterized by $\theta$. These parameters $\theta$ are obtained via optimizing a distance or discrepancy (or an approximation of it) between the posterior $p(x|y)$ and the $q_{\theta}(x)$.
Here, we focus on semi-implicit variational inference (SIVI) (Yin and Zhou, 2018). It enables a rich variational family by utilizing variational distributions, which we refer to as semi-implicit distributions (SID), defined as
$q_{k,r}(x)=\int k{}(x|z)r(z)\,\mathrm{d}z,$ (1)
where $k:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}_{+}$ is a kernel satisfying $\int k{}(x|z)\mathrm{d}x=1$; $r\in\mathcal{P}(\mathbb{R}^{d_{z}})$ is the mixing distribution and $\mathcal{P}(\mathbb{R}^{d_{z}})$ denotes the space of distributions with support $\mathbb{R}^{d_{z}}$ and finite second moments: $\mathbb{E}_{r}[\|z\|^{2}]<\infty$. SIDs are very flexible (Yin and Zhou, 2018) and can express complex properties, such as skewness, multimodality, and kurtosis, which typical variational families may fail to capture but are present in the posterior. There are various approaches to parameterizing these variational distributions: current techniques utilize neural networks built on top of existing kernels (e.g., Gaussian kernel) to define complex kernels (Titsias and Ruiz, 2019), and/or utilize pushforward distributions (a.k.a., implicit distributions) (Huszár, 2017; Yin and Zhou, 2018). On deciding a parameterization, an approximation to the posterior is obtained by minimizing the exclusive Kullback-Leibler (KL) divergence. This optimization has the same solution as minimizing the free energy (or the negative evidence lower bound) defined as
@@
However, since the integral in $q_{k,r}$ is typically intractable, directly optimizing $\mathcal{E}$ is not feasible. As a result, SIVI algorithms focus on designing tractable objectives by using upper bounds of $\mathcal{E}$ (Yin and Zhou, 2018); expensive Markov Chain Monte Carlo (MCMC) chains to estimate the gradient of $\mathcal{E}$ (Titsias and Ruiz, 2019); and score matching which results in mini-max objectives (Yu and Zhang, 2023).
In our work, we propose an alternative parameterization for SID: kernels are constructed as before (with parameter space denoted by $\Theta$) whereas the mixing distribution $r$ is obtained by optimizing over the whole space $\mathcal{P}(\mathbb{R}^{d_{z}})$. We motivate the case for minimizing a regularized version of the free energy $\mathcal{E}$ denoted by $\mathcal{E}_{\lambda}$ (see Eq. 4); thus, SIVI can be posed as the following optimization problem: $\operatorname*{arg\,min}_{(\theta,r)\in\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})}\mathcal{E}_{\lambda}(\theta,r)$. As a means to solving the SIVI problem, we construct a gradient flow that minimizes $\mathcal{E}_{\lambda}$ where the space $\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})$ is equipped with the Euclidean–Wasserstein geometry (Jordan et al., 1998; Kuntz et al., 2023). Via discretization, we obtain a practical algorithm for SIVI called Particle Variational Inference (PVI) that does not rely upon upper bounds of $\mathcal{E}$, MCMC chains, or solving minimax objectives.
@@
The structure of this paper is as follows: in Section 2, we begin with a discussion about prior approaches to parameterizing SIDs and their relation with one another. Then, in Section 3, we show how PVI is developed: beginning with designing a well-defined loss functional, construction of the gradient flow, and obtaining a practical algorithm. After, in Section 4, we study properties of a related gradient flow; and, in Section 5, conclude with experiments to demonstrate the efficacy of our proposal. Our related works are discussed in App. A.
2 On implicit mixing distributions in SID
@@
In this section, we outline prior approaches to parameterizing SIDs with implicit distributions and how that affects the resulting variational family. Before we begin, we shall outline key assumptions of SIVI. The kernel $k$ is assumed to be a reparametrized distribution in the sense of Salimans and Knowles (2013); Kingma and Welling (2014); Ruiz et al. (2016). In other words, the kernel $k$ is defined by the pair $(\phi,p_{k})$ where $\phi:\mathbb{R}^{d_{z}}\times\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ and $p_{k}\in\mathcal{P}(\mathbb{R}^{d_{x}})$ such that $\phi(z,\cdot)_{\#}p_{k}{=}k{}(\cdot|z).$ Furthermore, to ensure that it admits a tractable density, the map $\epsilon\mapsto\phi(z,\epsilon)$ is assumed to be a diffeomorphism for all $z\in\mathbb{R}^{d_{z}}$ with its inverse map is written as $\phi^{-1}(z,\cdot)$. From the change-of-variable formula, its density is given as $k{}(\cdot|z)=p_{k}(\phi^{-1}(z,\cdot))\,|\det\nabla_{x}\phi^{-1}(z,\cdot)|$. We sometimes write $k_{\phi,p_{k}}$ to denote the underlying $\phi$ and $p_{k}$ explicitly. Furthermore, the kernel $k$ is assumed to be computable and differentiable with respect to both arguments.
@@
Several approaches to the parameterization of SID have been explored in the literature. One can define variational families by choosing the kernel and mixing distribution from sets $\mathcal{K}$ and $\mathcal{R}$ respectively, i.e, the variational family is $\mathcal{Q}(\mathcal{K},\mathcal{R}):=\{q_{k,r}:k\in\mathcal{K},r\in\mathcal{R}\}$. Yin and Zhou (2018) focused on a fixed kernel $k$ with $r$ being a pushforward (or “implicit”) distribution, i.e., $r\in\{g_{\#}p_{r}:g\in\mathcal{G}\}=:\mathcal{R}_{\mathcal{G};p_{r}}$ where $\mathcal{G}$ is a subset of measurable mappings from the sample space of $p_{r}$ to $\mathbb{R}^{d_{z}}$. Thus, the $\mathcal{Q}_{\mathtt{YiZ}}$-variational family is $\mathcal{Q}(\{k\},\mathcal{R}_{\mathcal{G};p_{r}})$. On the other hand, Titsias and Ruiz (2019) considered a fixed mixing distribution $r$ with $k$ belonging to some parameterized class $\mathcal{K}$. The typical example is one where each kernel is defined by composing an existing kernel $k_{\phi,p_{k}}$ with a function $f\in\mathcal{F}$, the result is $k_{f;\phi,p_{k}}(\cdot|z):=k_{\phi(f(\cdot),\cdot),p_{k}}(\cdot|z)=\phi(f(z),\cdot)_{\#}p_{k}$ which clearly satisfies the reparameterization assumption. We denoted this kernel class as $\mathcal{K}_{\mathcal{F};\phi,p_{k}}:=\{k_{f;\phi,p_{k}}:f\in\mathcal{F}\}$ and its respective $\mathcal{Q}_{\mathtt{TR}}$-variational family is $\mathcal{Q}(\mathcal{K}_{\mathcal{G};\phi,p_{k}},\{r\})$. In Yu and Zhang (2023), they parameterized both $k$ and $r$, i.e., the $\mathcal{Q}_{\mathtt{YuZ}}$-variational family is $\mathcal{Q}(\mathcal{K}_{\mathcal{F};\phi,p_{k}},\mathcal{R}_{\mathcal{G};p_{r}})$. At first glance, it might seem that $\mathcal{Q}_{\mathtt{YuZ}}$ defines a larger variational family than either of the other approaches; however, under typical parameterizations, we show that they define the same variational family.
@@
Given a $\mathcal{Q}_{\mathtt{YuZ}}$-variational family, (i.e., $\mathcal{Q}_{\mathtt{YuZ}}:=\mathcal{Q}(\mathcal{K}_{\mathcal{F};\phi,p_{k}},\mathcal{R}_{\mathcal{G};p_{r}})$) then there is a $\mathcal{Q}_{\mathtt{YiZ}}$-variational family and $\mathcal{Q}_{\mathtt{TR}}$-variational family (i.e., $\mathcal{Q}_{\mathtt{TR}}:=\mathcal{Q}(\mathcal{K}_{\mathcal{F}\circ\mathcal{G};\phi,p_{k}},\{p_{r}\})$ and $\mathcal{Q}_{\mathtt{YiZ}}:=\mathcal{Q}({\{k_{\phi,p_{k}}}\},\mathcal{R}_{\mathcal{F}\circ\mathcal{G};p_{r}})$) such that $\mathcal{Q}_{\mathtt{YuZ}}=\mathcal{Q}_{\mathtt{YiZ}}=\mathcal{Q}_{\mathtt{TR}}$.
The proof can be found in App. D. This proposition shows that $\mathcal{Q}_{\mathtt{YuZ}}$-parameterizations defines the “same” variational faimily as $\mathcal{Q}_{\mathtt{YiZ}}$ and $\mathcal{Q}_{\mathtt{TR}}$. In practice, $\mathcal{F}$ and $\mathcal{G}$ are parameterized by neural networks hence $\mathcal{Q}_{\mathtt{YuZ}}$ can be viewed as $\mathcal{Q}_{\mathtt{YiZ}}$ or $\mathcal{Q}_{\mathtt{TR}}$ with a deeper neural networks $\mathcal{F}\circ\mathcal{G}$. This simplification is a direct result of using push-forward distributions. Although these models have shown promise Goodfellow et al. (2020), they have issues with expressivity particularly when distributions are disconnected Salmona et al. (2022). In our work, we utilize $\mathcal{Q}_{\mathtt{YuZ}}$-variational families, but, we instead propose to directly optimize on $\mathcal{P}(\mathbb{R}^{d_{z}})$ rather than using parametric implicit distributions; as a result, $\mathcal{Q}_{\mathtt{YuZ}}$ does not result simply reduce to $\mathcal{Q}_{\mathtt{YiZ}}$ or $\mathcal{Q}_{\mathtt{TR}}$.
@@ 3 Particle Variational Inference
In this section, we present our proposed method for SIVI, called particle variational inference (PVI). Similar to prior SIVI methods, the algorithm utilizes kernels (denoted by $k_{\theta}$) with parameters $\Theta$ which satisfy the assumptions listed in Section 2. One example is $k_{\theta}\in\mathcal{K}_{\Theta;\phi,p_{k}}$ where $\Theta$ is a function space parameterized by a neural network. We abuse the notation $\Theta$ to also indicate its corresponding weight space $\mathbb{R}^{d_{\theta}}$. The novelty of this algorithm is that, for the mixing distribution, we directly optimize over the space $\mathcal{P}(\mathbb{R}^{d_{z}})$ which loosens the requirement for the neural network in the kernel to learn complex mappings which results in a “simpler” optimization procedure and increases expressivity over prior methods. Thus, the variational parameters of SVI are $(\theta,r)\in\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})=:\mathcal{M}$ with its corresponding variational distribution defined as $q_{\theta,r}:=\int k_{\theta}(\cdot|z)r(z)\mathrm{d}z$. PVI arises naturally as a discretization of a gradient flow minimizing a suitably defined free energy on $\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})$ endowed with the Euclidean–Wasserstein geometry (Jordan et al., 1998; Ambrosio et al., 2005; Kuntz et al., 2023).

### v3

@@ Abstract
Semi-implicit variational inference (SIVI) enriches the expressiveness of variational families by utilizing a kernel and a mixing distribution to hierarchically define the variational distribution. Existing SIVI methods parameterize the mixing distribution using implicit distributions, leading to intractable variational densities. As a result, directly maximizing the evidence lower bound (ELBO) is not possible, so they resort to one of the following: optimizing bounds on the ELBO, employing costly inner-loop Markov chain Monte Carlo runs, or solving minimax objectives. In this paper, we propose a novel method for SIVI called Particle Variational Inference (PVI) which employs empirical measures to approximate the optimal mixing distributions characterized as the minimizer of a free energy functional. PVI arises naturally as a particle approximation of a Euclidean–Wasserstein gradient flow and, unlike prior works, it directly optimizes the ELBO whilst making no parametric assumption about the mixing distribution. Our empirical results demonstrate that PVI performs favourably compared to other SIVI methods across various tasks. Moreover, we provide a theoretical analysis of the behaviour of the gradient flow of a related free energy functional: establishing the existence and uniqueness of solutions as well as propagation of chaos results.
@@ 1 Introduction
In Bayesian inference, a quantity of vital importance is the posterior $p(x|y)={p(x,y)}/{\int p(x,y)\mathrm{d}x}$, where $p(x,y)$ is a probabilistic model, $y$ denotes the observed data, and $x$ the latent variable. An ever-present issue in Bayesian inference is that the posterior is often intractable. This is because the normalizing constant is available only in the form of an integral, and approximation methods are required. One popular method is variational inference (VI) (Jordan,, 1999; Wainwright and Jordan,, 2007; Blei et al.,, 2017). The essence of VI is to approximate the posterior with a member from a variational family $\mathcal{Q}$ where each element of $\mathcal{Q}$ is a distribution $q_{\theta}$ (called “variational distribution”) parameterized by $\theta$. These parameters $\theta$ are obtained via minimizing a distance or discrepancy (or an approximation of it) between the posterior $p(\cdot|y)$ and the variational distribution $q_{\theta}$.
Here, we focus on semi-implicit variational inference (SIVI) (Yin and Zhou,, 2018). It enables a rich variational family by utilizing variational distributions, which we refer to as semi-implicit distributions (SIDs), defined as
$q_{k,r}(x):=\int k{}(x|z)r(z)\,\mathrm{d}z,$ (1)
where $k:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}_{+}$ is a kernel satisfying $\int k{}(x|z)\mathrm{d}x=1$; $r\in\mathcal{P}(\mathbb{R}^{d_{z}})$ is the mixing distribution and $\mathcal{P}(\mathbb{R}^{d_{z}})$ denotes the space of distributions with support $\mathbb{R}^{d_{z}}$, with its usual Borel $\sigma$-field, with finite second moments. Here, and throughout, we assume that the distributions and kernels of interest admit densities. SIDs are very flexible (Yin and Zhou,, 2018) and can express complex properties, such as skewness, multimodality, and kurtosis. These properties might be present in the posterior but typical variational families may fail to capture them. There are various approaches to parameterizing these variational distributions: current techniques utilize neural networks built on top of existing kernels (e.g., Gaussian kernels) to define more complex kernels (Titsias and Ruiz,, 2019), and/or utilize pushforward distributions (a.k.a., implicit distributions (Huszár,, 2017)) (Yin and Zhou,, 2018). On choosing a parameterization, an approximation to the posterior is obtained by minimizing the exclusive Kullback-Leibler (KL) divergence. This optimization has the same solution as minimizing the free energy (or the negative evidence lower bound) defined as
@@
However, since the integral in $q_{k,r}$ is typically intractable, directly optimizing $\mathcal{E}$ is not feasible. As a result, SIVI algorithms focus on designing tractable objectives by using upper bounds of $\mathcal{E}$ (Yin and Zhou,, 2018); expensive Markov Chain Monte Carlo (MCMC) chains to estimate the gradient of $\mathcal{E}$ (Titsias and Ruiz,, 2019); and optimizing different objectives such as score matching which results in min-max objectives (Yu and Zhang,, 2023).
In our work, we propose an alternative parameterization for SIDs: kernels are constructed as before (with parameter space denoted by $\Theta$) whereas the mixing distribution $r$ is obtained by optimizing over the whole space $\mathcal{P}(\mathbb{R}^{d_{z}})$. We motivate the case for minimizing a regularized version of the free energy $\mathcal{E}$ denoted by $\mathcal{E}_{\lambda}$ (see Eq. 4); thus, SIVI can be posed as the following optimization problem: $\operatorname*{arg\,min}_{(\theta,r)\in\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})}\mathcal{E}_{\lambda}(\theta,r)$. As a means to solving the SIVI problem, we construct a gradient flow that minimizes $\mathcal{E}_{\lambda}$ where the space $\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})$ is equipped with the Euclidean–Wasserstein geometry (Jordan et al.,, 1998; Kuntz et al.,, 2023). Via discretization, we obtain a practical algorithm for SIVI called Particle Variational Inference (PVI) that does not rely upon upper bounds of $\mathcal{E}$, MCMC chains, or solving minimax objectives.
@@
The structure of this paper is as follows: in Section 2, we begin with a discussion of previous approaches to parameterizing SIDs and their relationship with one another. Then, in Section 3, we show how PVI is developed: beginning with designing a well-defined loss functional, the construction of the gradient flow, and how to obtain a practical algorithm. In Section 4, we study properties of a related gradient flow; and, in Section 5, we conclude with experiments to demonstrate the efficacy of our proposal. For sake of brevity, we defer our discussion of related works to App. A.
2 On implicit mixing distributions in SIDs
@@
This section outlines existing approaches to parameterizing SIDs with implicit distributions and how these choices affect the resulting variational family. Before we begin, we shall summarize the key assumptions of SIVI. The kernel $k$ is assumed to be a reparametrized distribution in the sense of Salimans and Knowles, (2013); Kingma and Welling, (2014); Ruiz et al., (2016). In other words, the kernel $k$ is defined by the pair $(\phi,p_{k})$ where $\phi:\mathbb{R}^{d_{z}}\times\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ and $p_{k}\in\mathcal{P}(\mathbb{R}^{d_{x}})$ such that $k{}(\cdot|z)=\phi(z,\cdot)_{\#}p_{k}$ Furthermore, to ensure that it admits a tractable density, the map $\epsilon\mapsto\phi(z,\epsilon)$ is assumed to be a diffeomorphism for all $z\in\mathbb{R}^{d_{z}}$ with its inverse map written as $\phi^{-1}(z,\cdot)$. From the change-of-variable formula, its density is given as $k{}(\cdot|z)=p_{k}(\phi^{-1}(z,\cdot))\,|\det\nabla_{x}\phi^{-1}(z,\cdot)|$. We sometimes write $k_{\phi,p_{k}}$ to denote the underlying $\phi$ and $p_{k}$ explicitly. Furthermore, the kernel $k$ is assumed to be computable and differentiable with respect to both arguments.
@@
Several approaches to the parameterization of SIDs have been explored in the literature. One can define the variational family by choosing the kernel and mixing distribution from sets $\mathcal{K}$ and $\mathcal{R}$ respectively, i.e., the variational family is $\mathcal{Q}(\mathcal{K},\mathcal{R}):=\{q_{k,r}:k\in\mathcal{K},r\in\mathcal{R}\}$. Yin and Zhou, (2018) focused on a fixed kernel $k$ with $r$ being a pushforward (or “implicit”) distribution, i.e., $r\in\{g_{\#}p_{r}:g\in\mathcal{G}\}=:\mathcal{R}_{\mathcal{G};p_{r}}$ where $\mathcal{G}$ is a subset of measurable mappings from the sample space of $p_{r}$ to $\mathbb{R}^{d_{z}}$. Thus, the $\mathcal{Q}_{\mathtt{YiZ}}$-variational family is $\mathcal{Q}(\{k\},\mathcal{R}_{\mathcal{G};p_{r}})$. On the other hand, Titsias and Ruiz, (2019) considered a fixed mixing distribution $r$ with $k$ belonging to some parameterized class $\mathcal{K}$. The typical example is one in which each kernel is defined by composing an existing kernel $k_{\phi,p_{k}}$ with a function $f\in\mathcal{F}$, the result is $k_{f;\phi,p_{k}}(\cdot|z):=k_{\phi(f(\cdot),\cdot),p_{k}}(\cdot|z)=\phi(f(z),\cdot)_{\#}p_{k}$ which clearly satisfies the reparameterization assumption. We denoted this kernel class as $\mathcal{K}_{\mathcal{F};\phi,p_{k}}:=\{k_{f;\phi,p_{k}}:f\in\mathcal{F}\}$ and its respective $\mathcal{Q}_{\mathtt{TR}}$-variational family is $\mathcal{Q}(\mathcal{K}_{\mathcal{G};\phi,p_{k}},\{r\})$. In Yu and Zhang, (2023), they combine both parameterization for $\mathcal{K}$ and $\mathcal{R}$, i.e., the $\mathcal{Q}_{\mathtt{YuZ}}$-variational family is $\mathcal{Q}(\mathcal{K}_{\mathcal{F};\phi,p_{k}},\mathcal{R}_{\mathcal{G};p_{r}})$. We note that this is how the variational family is presented in Yu and Zhang, (2023, see Sec. 2) but the authors used $\mathcal{Q}_{\texttt{TR}}$-variational family in experiments, i.e., $r$ was fixed. While $\mathcal{Q}_{\mathtt{YuZ}}$ might seem like it defines a larger variational family than the other approaches, under these common parameterization practices, we show that they define the same variational family.
@@
Given a $\mathcal{Q}_{\mathtt{YuZ}}$-variational family of the form $\mathcal{Q}_{\mathtt{YuZ}}:=\mathcal{Q}(\mathcal{K}_{\mathcal{F};\phi,p_{k}},\mathcal{R}_{\mathcal{G};p_{r}})$, then there is a $\mathcal{Q}_{\mathtt{YiZ}}$-variational family and $\mathcal{Q}_{\mathtt{TR}}$-variational family (i.e., $\mathcal{Q}_{\mathtt{TR}}:=\mathcal{Q}(\mathcal{K}_{\mathcal{F}\circ\mathcal{G};\phi,p_{k}},\{p_{r}\})$ and $\mathcal{Q}_{\mathtt{YiZ}}:=\mathcal{Q}({\{k_{\phi,p_{k}}}\},\mathcal{R}_{\mathcal{F}\circ\mathcal{G};p_{r}})$) such that $\mathcal{Q}_{\mathtt{YuZ}}=\mathcal{Q}_{\mathtt{YiZ}}=\mathcal{Q}_{\mathtt{TR}}$.
The proof can be found in App. D. This proposition shows that $\mathcal{Q}_{\mathtt{YuZ}}$-parameterization defines the “same” variational family as $\mathcal{Q}_{\mathtt{YiZ}}$ and $\mathcal{Q}_{\mathtt{TR}}$ when we parametrize $\mathcal{R}$ with push-forward distributions. In practice, $\mathcal{F}$ and $\mathcal{G}$ are parameterized by neural networks hence $\mathcal{Q}_{\mathtt{YuZ}}$ can be viewed as $\mathcal{Q}_{\mathtt{YiZ}}$ or $\mathcal{Q}_{\mathtt{TR}}$ with a deeper neural networks $\mathcal{F}\circ\mathcal{G}$. This simplification is a direct result of using push-forward distributions. Although this parametrization has shown promise e.g., Goodfellow et al., (2020)), they have issues with expressivity particularly when distributions are disconnected (Salmona et al.,, 2022). In our work, we follow in $\mathcal{Q}_{\mathtt{YuZ}}$-variational families, but, we avoid the use of push-forward distributions. Instead, we propose to directly optimize over $\mathcal{P}(\mathbb{R}^{d_{z}})$ and so, our variational family does not simply reduce to $\mathcal{Q}_{\mathtt{YiZ}}$ or $\mathcal{Q}_{\mathtt{TR}}$.
@@ 3 Particle Variational Inference
In this section, we present our proposed method for SIVI, called particle variational inference (PVI). Similar to prior SIVI methods, the algorithm utilizes kernels (denoted by $k_{\theta}$) with parameters $\Theta$ which satisfy the assumptions listed in Section 2. One example is $k_{\theta}\in\mathcal{K}_{\Theta;\phi,p_{k}}$ where $\Theta$ is a function space induced by a neural network. We slightly abuse the notation $\Theta$ to also indicate its corresponding weight space $\mathbb{R}^{d_{\theta}}$. The novelty of this algorithm is that, for the mixing distribution, we directly optimize over the space $\mathcal{P}(\mathbb{R}^{d_{z}})$ which loosens the requirement for the neural network in the kernel to learn complex mappings. The result is a “simpler” optimization procedure and increases expressivity over existing methods. Thus, the variational parameters of PVI are $(\theta,r)\in\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})=:\mathcal{M}$ with its corresponding variational distribution defined as $q_{\theta,r}:=\int k_{\theta}(\cdot|z)r(z)\mathrm{d}z$. PVI arises naturally as a discretization of a gradient flow minimizing a suitably defined free energy on $\Theta\times\mathcal{P}(\mathbb{R}^{d_{z}})$ endowed with the Euclidean–Wasserstein geometry (Jordan et al.,, 1998; Ambrosio et al.,, 2005; Kuntz et al.,, 2023). In Section 3.1, we begin by constructing a suitably defined free energy functional; then, in Section 3.2, we formulate its gradient flow; finally, in Section 3.3, we construct PVI from its gradient flow.